OpenStudy (anonymous):

please answer with solution :) find the length of the hypotenuse of an isosceles right triangle whose area is 18 square units

OpenStudy (anonymous):

answer6

OpenStudy (dls):

Area is given by:  So..36=b*h.. but b=h,since its isoceles,so.. 36=b^2 b=6 Now use Pythagoras